Assume that adults have IQ scores that are normally distributed with a mean of 104.3 and a standard deviation of 22.8. Find the probability that a randomly selected adult has an IQ greater than 147.2.

​(Hint: Draw a​ graph.)
Answer #1

Solution :

Given that,

mean = = 104.3

standard deviation = = 22.8

P(x >147.2 ) = 1 - P(x<147.2 )

= 1 - P[(x -) / < (147.2 -104.3) /22.8 ]

= 1 - P(z <1.88 )

Using z table

= 1 -  0.9699

probability= 0.0301